Mickey is 5 years older than Minnie. Three more than four times Minnie's age is 2 less than 3 times Mickey's age. How old is each person?
Answer by stanbon(75887)

Mickey is 5 years older than Minnie. Three more than four times Minnie's age is 2 less than 3 times Mickey's age. How old is each person?
----
Equations:
k = n + 5
4n+3 = 3k-2
------------
Substitute for "k" and solve for "n":
4n+3 = 3(n+5)-2
4n+3 = 3n+13
n = 10 (Minnie's age now)
---
solve for "k":
k = n + 5
k = 15 (Mickey's age now)
===============================
Cheers,
Stan H.
===============